OpenStudy (saifoo.khan):

Differentiate: v = 6t + 4 cos 2t.

OpenStudy (amistre64):

piece of cake

OpenStudy (anonymous):

\[6-8sin(2t)\]

OpenStudy (amistre64):

v' = 6 -8cos(2t)

OpenStudy (saifoo.khan):

help me eating it/ ;)

OpenStudy (anonymous):

6 - 8sin2t

OpenStudy (saifoo.khan):

@amistre: it will b sin*

OpenStudy (amistre64):

did I not type a sin? lol

OpenStudy (anonymous):

v' = 6 - 8 sin2t since d/dx sin2t = -2cos2t

OpenStudy (anonymous):

o whoops

OpenStudy (anonymous):

i mean d/dx cos = -sin

OpenStudy (amistre64):

my head said sin; and my fingers decide to go another route

OpenStudy (saifoo.khan):

can anybdy help me tht where can i find these formulaes?? :/

OpenStudy (saifoo.khan):

lol. @amisrte

OpenStudy (amistre64):

the trig derives are as much memorization as the trig functions themselves

OpenStudy (amistre64):

but; we can try to do it thru th elimit definition

OpenStudy (saifoo.khan):

where i can find its formulaes?

OpenStudy (saifoo.khan):

all of them.

OpenStudy (amistre64):

\[\lim_{h->0}\frac{cos(x+h)-cos(x)}{x+h-x}\]

OpenStudy (amistre64):

in the back or front of any calculus textbook id assume

OpenStudy (amistre64):

\[cos(x)cos(h)-sin(x)sin(h)-cos(x)//h\] \[cos(x)(cos(h)-1)-sin(x)sin(h)//h\]

OpenStudy (amistre64):

\[\lim_{h->0} cos(x)\frac{cos(h-1)}{h}-\lim_{h->0}sin(x)\frac{sin(h)}{h}\]

OpenStudy (amistre64):

\[{cos(h)\over h} ->0\] \[\frac{sin(h)}{h}->1\] leaning us with: \[-sin(x)\]

OpenStudy (saifoo.khan):

k, thnks a TON! :D

OpenStudy (amistre64):

had a typo up in there, but im sure you can see thru it ;)

OpenStudy (amistre64):

rule of thumb; if it starts with a "c"; the derivative is "-" cos = -sin cot = -csc^2 csc = -csc cot

OpenStudy (amistre64):

the others or just the evil twins .. so to speak sin = cos tan = sec^2 sec = sec tan

OpenStudy (saifoo.khan):

k, thnks a TON! :D

OpenStudy (saifoo.khan):

heyyy! plz show me the working oif: 4 cos 2t

OpenStudy (saifoo.khan):

how u got: 8sin(2t)

OpenStudy (amistre64):

well, it amounts to the 'chain rule' and the 'product rule' if you know the shortcuts. (rl)' = r'l +rl' and the chain rule is: F[g(x)] = F'[g(x)] * g'(x)

OpenStudy (amistre64):

r=4 ; l=cos(2x) , F(g(x))=cos(2x) , g(x) = 2x r'=0 ; l'= -sin(2x) * 2 [4 cos(2x)]' = 0(cos(2x)) +4(-2 sin(2x)) = -8sin(2x)

OpenStudy (amistre64):

it looks more complicated than it is; some things just dont notate right on screen